To make 1.0 L of buffered solution of a 10.0 pH and a total [ ] of carbonate and hydrocarbonate is 0.10 M, how many grams of Na2CO3 and NaHCO3 are needed?

Carbonic acid Ka1 = 4.3 * 10-7 and Ka2 = 5.6 * 10-11
Answer #1

We have, pKa = - log Ka

\therefore pKa = - log ( 5.6 \times 10 -11) = 10.25

pH of buffer solution is calculated by using Henderson's equation.

pH = pKa + log [ Na2CO3 ] / [ NaHCO3 ]

10.0 = 10.25 + log [ Na2CO3 ] / [ NaHCO3 ]

log [ Na2CO3 ] / [ NaHCO3 ] = 10.0 - 10.25 = - 0.25

[ Na2CO3 ] / [ NaHCO3 ] = 10 -0.25 = 0.562

[ Na2CO3 ] / [ NaHCO3 ] = 0.562

i e [ Na2CO3 ] = 0.562 [ NaHCO3 ] ---------------------->(1)

We have, [ Na2CO3 ] + [ NaHCO3 ] = 0.10 M

From equation 1 , we can write

0.562 [ NaHCO3 ] + [ NaHCO3 ] = 0.10

1.562 [ NaHCO3 ] = 0.10

[ NaHCO3 ] = 0.10 / 1.562

[ NaHCO3 ] =0.0640 M

\therefore [ Na2CO3 ] = 0.562 [ NaHCO3 ] becomes [ Na2CO3 ] = 0.562 ( 0.0640 ) = 0.0360 M

Volume of buffer solution = 1.0 L

We know that, Molarity = No. of moles of Solute / volume of solution in L

\therefore [ NaHCO3 ] = No. of moles of  NaHCO3 / volume of solution in L

\therefore 0.0640 mol / L = No. of moles of  NaHCO3 / 1.0 L

No. of moles of  NaHCO3 = 0.0640 mol / L \times 1.0 L = 0.0640 mol

No. of moles = Mass / Molar mass

\therefore Mass of NaHCO3 = 0.0640 mol ( 84.00 g/ mol ) = 5.376 g

Calculation for Na2CO3

[ Na2CO3 ] =No. of moles of Na2CO3 / volume of solution in L

\therefore 0.0360 mol / L = No. of moles of Na2CO3 / 1.0 L

No. of moles of Na2CO3 = 0.0360 mol / L \times 1.0 L = 0.0360 mol

We know that, No. of moles = Mass / Molar mass

\therefore Mass of Na2CO3= 0.0360 mol ( 105.99 g/ mol ) = 3.816 g

ANSWER : Mass of NaHCO3 = 5.376 g

Mass of Na2CO3= 3.816 g